An international long distance phone plan has a $5.99 monthly fee plus $.59 per minute for all calls to Canada. Express the formula using the function notation C(m) as the monthly charge for m minutes of long distance calls to Canada.
The function is C(m)= ___
The charge for 95 minutes of long distance calls to Canada is C (95)= $___
The charge for 159 minutes of long distance calls to Canada is C (159) = $___

Answer by CubeyThePenguin(3113)

C(m) = 5.99 + 0.59m

C(95) = 5.99 + 0.59(95) = $62.04

C(159) = 5.99 + 0.59(159) = $99.80